# via Irnerio 46, I-40126 Bologna, Italy

### v2

@@ 2.2 Feynman integral families
In figs.˜1, 2 and 3 we depict the graphs associated with the top sectors of the families, indicating the defining set of exponents and the superfamily. To determine the MIs of each family, we generate the IBP relations using FFIntRed [^1] An in-house Mathematica package by Tiziano Peraro. and we solve the resulting systems with FiniteFlow Peraro:2019svx. The number of MIs obtained for each family, in both the equal- and different-mass case, is summarized in table˜1. The families in the different-mass case contain noticeably more MIs than those in the equal-mass case, with the maximum numbers being 344 and 196, respectively. Interestingly, while $\text{PT}_{1}$ is the family with the highest number of MIs in the different-mass case, this is no longer true for the equal-mass configuration, where $\text{PT}_{2}$ becomes the largest one. This behavior is expected, as $\text{PT}_{1}$ becomes topologically more symmetric in the equal-mass limit, leading to a reduction in the number of independent MIs. Finally, we note that the number of top-sector MIs is generally identical between the two mass configurations, except for $\text{PT}_{1}$, which contains one fewer MI in the equal-mass case.
@@ 3.1 Construction of the pure bases
A crucial step is the selection of the pre-canonical candidates, for which we employed several complementary techniques. For sectors with only one or a few MIs, educated guesses involving squared propagators in lower sectors and ISPs in higher sectors were typically sufficient. For more complicated cases, we applied the methods of building blocks Wasser:2018qvj and integrand analysis Arkani-Hamed:2010pyv; Henn:2020lye; Wasser:2022kwg. For the latter, we worked both in four dimensions using the package DlogBasis Henn:2020lye, including some tricks to handle massive external momenta Papadopoulos:2014lla; Canko:2021hvh, and in $d$ dimensions using the Baikov representation Baikov:1996iu; Baikov:1996rk, for which we used the package Baikov.m Frellesvig:2017aai. After obtaining an $\varepsilon$-factorised form of the DEs on the maximal cut, we release the cuts and we bring the DEs into canonical form also in the sub-sectors. In practice, we found that most entries were already either $\varepsilon$-factorised or linear in $\varepsilon$. In the latter case, we achieved the canonical form by rotating the basis, analogously to what we did on the maximal cut.
In our study, we first derived the canonical bases of the families in the general kinematic configuration. We then obtained the bases for the equal-mass case by setting $m_{3}=m_{4}=m$ in the normalisation of our pure integrals and reducing them via IBPs in the equal-mass kinematics. The resulting bases were already in canonical form, requiring no further manipulation. For both mass configurations, the pure bases and the canonical DEs are provided in the ancillary files canko_2025_17727482.
@@ 3.2 Construction of the alphabet
There are several packages that implement algorithms to generate candidates for algebraic letters, such as BaikovLetter Jiang:2024eaj, which we employed in Canko:2024ara to identify the letters involving $r_{1}$. However, using BaikovLetter for the letters depending on $r_{2}$, $r_{4}$ and $r_{5}$ proved to be too time-consuming. For these cases, we instead used the package SOFIA Correia:2025yao, which automates the detection of the singularities of the Feynman integrals and constructs candidate letters utilizing Effortless. With the resulting ansatz for the alphabet, we were able to determine all entries of the connection matrices except for a few, involving $r_{1}$ and one of the new square roots. We constructed the missing letters following the same strategy of Effortless, described in Matijasic:2024too. Assuming that $q_{j}(\vec{s})=1$ in eq.˜23, we know that the product of numerator and denominator of each letter must factorise in the alphabet Heller:2019gkq; Zoia:2021zmb:

### v3

@@ 2.2 Feynman integral families
In figs.˜1, 2 and 3 we depict the graphs associated with the top sectors of the families, indicating the defining set of exponents and the superfamily. To determine the MIs of each family, we generate the IBP relations using FFIntRed [^1] An in-house Mathematica package by Tiziano Peraro. and we solve the resulting systems with FiniteFlow Peraro:2019svx . The number of MIs obtained for each family, in both the equal-mass and different-mass case, is summarized in table˜1. Across all families, we find 823 independent master integrals in the general case, and 523 in the equal-mass case. The families in the different-mass case contain noticeably more MIs than those in the equal-mass case, with the maximum numbers being 344 and 196, respectively. For comparison, the planar three-loop four-point one-mass families DiVita:2014pza ; Canko:2020gqp ; Canko:2021xmn ; Syrrakos:2023mor range from 83 to 166 MIs, while the non-planar case Henn:2023vbd ; Gehrmann:2024tds ; Henn:2024pki ; Guan:2025awp is comparable to our calculation, with families having from 114 up to 371 MIs, although the largest top-sector contains in that case 19 MIs. Interestingly, while $\text{PT}_{1}$ is the family with the highest number of MIs in the different-mass case, this is no longer true for the equal-mass configuration, where $\text{PT}_{2}$ becomes the largest one. This behavior is expected, as $\text{PT}_{1}$ becomes topologically more symmetric in the equal-mass limit, leading to a reduction in the number of independent MIs. Finally, we note that the number of top-sector MIs is generally identical between the two mass configurations, except for $\text{PT}_{1}$, which contains one fewer MI in the equal-mass case.
We verified, by means of IBP reductions, that all remaining planar three-loop topologies with two off-shell legs do not give rise to independent master integrals. Instead, they reduce either to master integrals belonging to the families considered in this work, or to products of lower-loop integrals, namely one-loop times two-loop integrals or products of three one-loop integrals. This confirms that the set of integral families studied here is complete for describing all the three-loop planar Feynman integrals with two off-shell legs and massless internal propagators.
@@ 3.1 Construction of the pure bases
A crucial step is the selection of the pre-canonical candidates, for which we employed several complementary techniques. For sectors with only one or a few MIs, educated guesses involving squared propagators in lower sectors and ISPs in higher sectors were typically sufficient. For more complicated cases, we applied the methods of building blocks Wasser:2018qvj and integrand analysis Arkani-Hamed:2010pyv ; Henn:2020lye ; Wasser:2022kwg . For the latter, we worked both in four dimensions using the package DlogBasis Henn:2020lye , including some tricks to handle massive external momenta Papadopoulos:2014lla ; Canko:2021hvh , and in $d$ dimensions using the Baikov representation Baikov:1996iu ; Baikov:1996rk , for which we used the package Baikov.m Frellesvig:2017aai . After obtaining an $\varepsilon$-factorised form of the DEs on the maximal cut, we release the cuts and we bring the DEs into canonical form also in the sub-sectors. In practice, we found that most entries were already either $\varepsilon$-factorised or linear in $\varepsilon$. In the latter case, we achieved the canonical form by rotating the basis, analogously to what we did on the maximal cut. In a few cases, notably the ten- and nine-propagators sectors of $\mathrm{PT}_{1}$ and $\mathrm{PT}_{2}$, the couplings to some of the lower sectors still involved factorised poles in $\varepsilon$ after releasing the cuts. This was a consequence of missing contributions to the canonical integrals coming from some reducible sub-sector, which we managed to find using DlogBasis.
It is known that the computational cost of reconstructing the DEs depends on their complexity, for example the degree of the rational functions involved. As the reconstruction of the canonical DEs [^2] Here we mean canonical up to a rotation by a diagonal matrix involving the square roots of eqs. 24 and 25. for $\mathrm{PT}_{1}$ takes $\mathcal{O}$(hours) on a cluster, it is clear that reconstructing the DEs in a non-canonical form would be time-consuming and inefficient. Fortunately, most of the steps in the procedure only require the knowledge of the $\varepsilon$-dependence of the differential equation, while the full kinematic dependence is only needed to determine the rotation from the pre-canonical candidates to the canonical basis. Therefore, in order to be more efficient and to be able to test many candidates, we resorted to numerical slices, reconstructing only the dependence of the DEs on $\varepsilon$.
In our study, we first derived the canonical bases of the families in the general kinematic configuration. We then obtained the bases for the equal-mass case by setting $m_{3}=m_{4}=m$ in the normalisation of our pure integrals and reducing them via IBPs in the equal-mass kinematics. The resulting bases were already in canonical form, requiring no further manipulation. For both mass configurations, the pure bases and the canonical DEs are provided in the ancillary files canko_2025_17727482 . Moreover, we refer to appendix˜B for the graphs of the genuinely new integrals appearing in our calculation.
@@ 3.2 Construction of the alphabet
There are several packages that implement algorithms to generate candidates for algebraic letters, such as BaikovLetter Jiang:2024eaj , which we employed in Canko:2024ara to identify the letters involving $r_{1}$. However, using BaikovLetter for the letters depending on $r_{2}$, $r_{4}$ and $r_{5}$ proved to be too time-consuming. For these cases, we instead used the package SOFIA Correia:2025yao , which automates the detection of the singularities of the Feynman integrals using the method proposed in Caron-Huot:2024brh and constructs candidate letters utilizing Effortless. With the resulting ansatz for the alphabet, we were able to determine all entries of the connection matrices except for a few, involving $r_{1}$ and one of the new square roots. We constructed the missing letters following the same strategy of Effortless, described in Matijasic:2024too . Assuming that $q_{j}(\vec{s})=1$ in eq.˜23, we know that the product of numerator and denominator of each letter must factorise in the alphabet Heller:2019gkq ; Zoia:2021zmb :
@@ Appendix A Ancillary files
In the folder info, we provide the following files:
  • MI_relations.m (MI_relations_same_mass.m): the relations between the canonical master integrals of different families;
  • independent_integrals.m (independent_integrals_same_mass.m): the list of 823 (523) independent master integrals across all families in the different-mass (equal-mass) case.
In the above files, the notation ut[fam,j] refers to the $j$-th entry in the list of master integrals of family fam. Moreover, the folder also contains the file iterated_integrals.wl, which generates the expressions of the MIs in terms of Chen iterated integrals Chen:1977oja . These are defined as
$I_{\gamma}\left(\omega_{1},...,\omega_{n};\lambda\right)=\int\limits_{a}^{\lambda}d\lambda_{1}\mathfrak{f}_{1}\left(\lambda_{1}\right)\int\limits_{a}^{\lambda_{1}}d\lambda_{2}\mathfrak{f}_{2}\left(\lambda_{2}\right)\dots\int\limits_{a}^{\lambda_{n-1}}d\lambda_{n}\mathfrak{f}_{n}\left(\lambda_{n}\right),$ (33)
where $\mathfrak{f}_{j}\left(\lambda\right)d\lambda=\gamma^{\ast}\omega_{j}$ is the pull-back of the one-form $\omega_{j}$ along a path $\gamma$. In the dlog case $\omega_{i}=\mathrm{d}\log\left(w_{k_{i}}\right)$, hence we use the short-hand notation
$\mathtt{II}[\mathtt{W[k_{1}]},\dots,\mathtt{W[k_{n}]}]:=I_{\gamma}\left(\omega_{1},...,\omega_{n};\lambda\right),$ (34)
where $k_{i}$ is the letter index.
@@
Appendix B New topologies
In this appendix, we collect the pure candidates for the genuinely new sectors in the general-mass case, referring to their position within the basis of the corresponding integral family (see the ancillary files canko_2025_17727482 ). We exclude the sectors presented in Appendix B of Ref. Canko:2024ara . Regarding the sector definition, we use the notation
$\text{SF}[\nu_{1},\ldots,\nu_{15}],$
where $\text{SF}=\{\text{$F_{123}$},\text{$F_{132}$}\}$ denotes the superfamily, and $\nu_{i}\in\{0,1\}$ are the exponents of the propagators of the corresponding superfamily. Pure candidates are denoted by
$I_{\text{pos}}^{\text{Fam}},$
where $\text{Fam}=\{\text{PL2},\text{PL3},\text{PT1},\text{PT2},\text{PT3}\}$ labels the integral family, and $\text{pos}\in\{1,\ldots,344\}$ specifies the position of the candidate within the corresponding basis. In total, we identified 76 new sectors, comprising 212 MIs. In the following, we classify the sectors according to the number of propagators they contain, starting with those with six propagators and ending with those with ten.
Six-Propagator Pure Candidates
Sector $\mathbf{F_{132}}$[0,1,0,0,1,0,0,0,0,0,1,1,1,0,1]
$\{I_{113}^{\text{PL3}}\}$
Sector $\mathbf{F_{132}}$[0,1,0,0,0,0,1,0,1,0,0,1,1,0,1]
$\{I_{79}^{\text{PL3}},I_{80}^{\text{PL3}},I_{81}^{\text{PL3}},I_{82}^{\text{PL3}},I_{83}^{\text{PL3}},I_{84}^{\text{PL3}}\}$
Sector $\mathbf{F_{132}}$[1,1,0,0,0,0,0,0,0,0,1,1,1,0,1]
$\{I_{97}^{\text{PL3}}\}$
Sector $\mathbf{F_{132}}$[0,1,0,0,0,0,1,1,1,0,0,0,1,0,1]
$\{I_{200}^{\text{PT3}}\}$
Sector $\mathbf{F_{132}}$[0,1,0,0,0,0,1,0,1,0,0,1,0,1,1]
$\{I_{171}^{\text{PT3}},I_{172}^{\text{PT3}}\}$
Sector $\mathbf{F_{123}}$[0,0,1,0,0,0,1,0,1,0,0,1,0,1,1]
$\{I_{281}^{\text{PT1}},I_{282}^{\text{PT1}}\}$
Sector $\mathbf{F_{123}}$[1,0,0,0,0,1,0,1,0,0,1,0,1,0,1]
$\{I_{86}^{\text{PL2}},I_{87}^{\text{PL2}}\}$
Sector $\mathbf{F_{132}}$[0,1,0,0,1,0,1,0,0,0,0,1,1,0,1]
$\{I_{94}^{\text{PL3}},I_{93}^{\text{PL3}}\}$
Seven-Propagator Pure Candidates
Sector $\mathbf{F_{123}}$[1,0,0,0,0,0,0,1,0,1,1,1,1,0,1]
$\{I_{65}^{\text{PL2}}\}$
Sector $\mathbf{F_{132}}$[0,1,0,0,0,0,1,0,1,0,1,1,1,0,1]
$\{I_{66}^{\text{PL3}}\}$
Sector $\mathbf{F_{123}}$[0,1,1,1,1,0,0,0,0,1,0,0,1,0,1]
$\{I_{129}^{\text{PT2}}\}$
Sector $\mathbf{F_{132}}$[1,1,1,0,0,0,1,0,0,0,0,1,0,1,1]
$\{I_{111}^{\text{PT3}}\}$
Sector $\mathbf{F_{123}}$[1,1,1,0,0,0,0,1,1,0,0,0,1,0,1]
$\{I_{161}^{\text{PT1}},I_{162}^{\text{PT1}}\}$
Sector $\mathbf{F_{123}}$[0,1,1,1,0,0,0,1,1,0,0,0,0,1,1]
$\{I_{108}^{\text{PT2}},I_{109}^{\text{PT2}}\}$
Sector $\mathbf{F_{132}}$[1,1,1,0,0,0,0,1,1,0,0,0,1,0,1]
$\{I_{105}^{\text{PT3}},I_{106}^{\text{PT3}}\}$
Sector $\mathbf{F_{123}}$[0,1,0,1,1,0,0,1,0,1,0,0,1,0,1]
$\{I_{132}^{\text{PT2}}\}$
Sector $\mathbf{F_{123}}$[0,0,1,0,0,0,1,1,1,0,0,1,0,1,1]
$\{I_{198}^{\text{PT1}}\}$
Sector $\mathbf{F_{132}}$[0,1,0,0,0,0,1,1,1,0,0,1,1,0,1]
$\{I_{130}^{\text{PT3}}\}$
Sector $\mathbf{F_{132}}$[1,1,0,0,0,0,1,1,0,0,0,1,1,1,0]
$\{I_{122}^{\text{PT3}}\}$
Sector $\mathbf{F_{123}}$[0,1,0,0,0,0,1,1,1,0,0,1,0,1,1]
$\{I_{177}^{\text{PT1}},I_{178}^{\text{PT1}}\}$
Sector $\mathbf{F_{123}}$[0,1,1,0,0,0,1,0,1,0,0,1,0,1,1]
$\{I_{171}^{\text{PT1}},I_{172}^{\text{PT1}}\}$
Sector $\mathbf{F_{123}}$[0,1,1,0,0,0,0,1,1,0,0,1,1,1,0]
$\{I_{153}^{\text{PT1}},I_{154}^{\text{PT1}},I_{155}^{\text{PT1}}\}$
Sector $\mathbf{F_{132}}$[0,1,1,0,0,0,0,1,1,0,0,1,1,1,0]
$\{I_{74}^{\text{PT3}},I_{75}^{\text{PT3}},I_{76}^{\text{PT3}},I_{77}^{\text{PT3}},I_{78}^{\text{PT3}},I_{79}^{\text{PT3}}\}$
Sector $\mathbf{F_{123}}$[1,0,0,0,0,1,0,1,0,0,1,1,1,0,1]
$\{I_{54}^{\text{PL2}},I_{55}^{\text{PL2}}\}$
Sector $\mathbf{F_{132}}$[0,1,0,0,1,0,1,0,1,0,0,1,1,0,1]
$\{I_{55}^{\text{PL3}},I_{56}^{\text{PL3}}\}$
Sector $\mathbf{F_{123}}$[1,1,1,0,0,0,0,1,0,0,0,1,1,1,0]
$\{I_{184}^{\text{PT1}}\}$
Sector $\mathbf{F_{123}}$[0,1,1,1,1,0,0,0,1,0,0,0,1,1,0]
$\{I_{128}^{\text{PT2}}\}$
Sector $\mathbf{F_{132}}$[1,1,1,0,0,0,0,1,0,0,0,1,1,1,0]
$\{I_{118}^{\text{PT3}}\}$
Sector $\mathbf{F_{132}}$[0,1,0,0,0,0,1,1,1,0,0,0,1,1,1]
$\{I_{113}^{\text{PT3}},I_{114}^{\text{PT3}}\}$
Sector $\mathbf{F_{132}}$[0,1,1,0,0,0,0,1,1,0,0,0,1,1,1]
$\{I_{61}^{\text{PT3}},I_{62}^{\text{PT3}},I_{63}^{\text{PT3}},I_{64}^{\text{PT3}},I_{65}^{\text{PT3}},I_{66}^{\text{PT3}},I_{67}^{\text{PT3}}\}$
Sector $\mathbf{F_{123}}$[1,0,0,1,0,0,0,1,0,1,1,0,1,0,1]
$\{I_{52}^{\text{PL2}},I_{53}^{\text{PL2}}\}$
Sector $\mathbf{F_{132}}$[1,0,1,0,0,0,0,1,1,0,0,1,1,0,1]
$\{I_{111}^{\text{PT3}},I_{112}^{\text{PT3}}\}$
Sector $\mathbf{F_{132}}$[0,1,1,0,0,0,1,0,1,0,0,1,1,0,1]
$\{I_{44}^{\text{PL3}},I_{45}^{\text{PL3}},I_{46}^{\text{PL3}},I_{47}^{\text{PL3}},I_{48}^{\text{PL3}}\}$
Sector $\mathbf{F_{132}}$[0,1,1,0,1,0,0,0,0,0,1,1,1,0,1]
$\{I_{34}^{\text{PL3}},I_{35}^{\text{PL3}},I_{36}^{\text{PL3}},I_{37}^{\text{PL3}},I_{38}^{\text{PL3}}\}$
Sector $\mathbf{F_{123}}$[0,1,0,0,0,0,1,1,1,0,0,1,1,1,0]
$\begin{split}\{&I_{90}^{\text{PT1}},I_{91}^{\text{PT1}},I_{92}^{\text{PT1}},I_{93}^{\text{PT1}},I_{94}^{\text{PT1}},\\ &I_{95}^{\text{PT1}},I_{96}^{\text{PT1}},I_{97}^{\text{PT1}},I_{98}^{\text{PT1}},I_{99}^{\text{PT1}},\\ &I_{100}^{\text{PT1}},I_{101}^{\text{PT1}},I_{102}^{\text{PT1}},I_{103}^{\text{PT1}},I_{104}^{\text{PT1}}\}\end{split}$
Eight-Propagator Pure Candidates
Sector $\mathbf{F_{123}}$[1,0,0,1,0,1,0,1,0,0,1,1,1,0,1]
$\begin{split}\{&I_{28}^{\text{PL2}}\}\end{split}$
Sector $\mathbf{F_{132}}$[0,1,1,0,1,0,1,0,1,0,0,1,1,0,1]
$\begin{split}\{&I_{28}^{\text{PL3}}\}\end{split}$
Sector $\mathbf{F_{123}}$[0,1,1,0,0,0,1,1,1,0,0,1,0,1,1]
$\{I_{89}^{\text{PT1}}\}$
Sector $\mathbf{F_{132}}$[0,1,1,0,0,0,1,1,1,0,0,1,1,0,1]
$\{I_{53}^{\text{PT3}}\}$
Sector $\mathbf{F_{132}}$[1,1,0,0,0,0,1,1,1,0,0,1,1,0,1]
$\{I_{51}^{\text{PT3}}\}$
Sector $\mathbf{F_{123}}$[0,1,1,0,0,0,1,1,1,0,0,1,1,1,0]
$\{I_{77}^{\text{PT1}},I_{78}^{\text{PT1}},I_{79}^{\text{PT1}}\}$
Sector $\mathbf{F_{123}}$[1,1,1,0,0,0,0,1,1,0,0,1,1,1,0]
$\begin{split}\{&I_{87}^{\text{PT1}}\}\end{split}$
Sector $\mathbf{F_{123}}$[0,1,1,1,1,0,0,1,1,0,0,0,1,1,0]
$\{I_{54}^{\text{PT2}}\}$
Sector $\mathbf{F_{132}}$[1,1,1,0,0,0,0,1,1,0,0,1,1,1,0]
$\{I_{50}^{\text{PT3}}\}$
Sector $\mathbf{F_{123}}$[1,0,0,0,0,1,0,1,0,1,1,1,1,0,1]
$\{I_{26}^{\text{PL2}},I_{27}^{\text{PL2}}\}$
Sector $\mathbf{F_{132}}$[0,1,0,0,1,0,1,0,1,0,1,1,1,0,1]
$\{I_{26}^{\text{PL3}},I_{27}^{\text{PL3}}\}$
Sector $\mathbf{F_{123}}$[0,1,1,1,0,0,0,1,1,1,0,0,0,1,1]
$\{I_{50}^{\text{PT2}},I_{51}^{\text{PT2}}\}$
Sector $\mathbf{F_{132}}$[1,1,1,0,0,0,1,1,1,0,0,0,1,0,1]
$\{I_{46}^{\text{PT3}},I_{47}^{\text{PT3}}\}$
Sector $\mathbf{F_{123}}$[0,1,1,1,1,0,0,1,0,1,0,0,1,0,1]
$\{I_{40}^{\text{PT2}},I_{41}^{\text{PT2}},I_{42}^{\text{PT2}}\}$
Sector $\mathbf{F_{132}}$[1,1,1,0,0,0,1,0,1,0,0,1,0,1,1]
$\{I_{41}^{\text{PT3}},I_{42}^{\text{PT3}},I_{43}^{\text{PT3}}\}$
Sector $\mathbf{F_{123}}$[1,1,1,0,0,0,1,0,1,0,0,1,0,1,1]
$\{I_{70}^{\text{PT1}},I_{71}^{\text{PT1}},I_{72}^{\text{PT1}},I_{73}^{\text{PT1}}\}$
Sector $\mathbf{F_{123}}$[1,1,0,1,0,0,0,0,0,1,1,1,1,0,1]
$\{I_{22}^{\text{PL2}},I_{23}^{\text{PL2}}\}$
Sector $\mathbf{F_{132}}$[1,1,1,0,0,0,0,0,1,0,1,1,1,0,1]
$\{I_{22}^{\text{PL3}},I_{23}^{\text{PL3}}\}$
Sector $\mathbf{F_{123}}$[1,1,0,0,0,0,0,1,0,1,1,1,1,0,1]
$\{I_{24}^{\text{PL2}},I_{25}^{\text{PL2}}\}$
Sector $\mathbf{F_{132}}$[1,1,0,0,0,0,1,0,1,0,1,1,1,0,1]
$\{I_{24}^{\text{PL3}},I_{25}^{\text{PL3}}\}$
Sector $\mathbf{F_{132}}$[0,1,1,0,1,0,0,0,1,0,1,1,1,0,1]
$\{I_{14}^{\text{PL3}},I_{15}^{\text{PL3}},I_{16}^{\text{PL3}},I_{17}^{\text{PL3}}\}$
Sector $\mathbf{F_{123}}$[1,0,0,1,0,1,0,0,0,1,1,1,1,0,1]
$\{I_{14}^{\text{PL2}},I_{15}^{\text{PL2}},I_{16}^{\text{PL2}},I_{17}^{\text{PL2}}\}$
Sector $\mathbf{F_{123}}$[1,1,1,0,0,0,0,1,1,0,0,1,1,0,1]
$\{I_{82}^{\text{PT1}},I_{83}^{\text{PT1}}\}$
Sector $\mathbf{F_{123}}$[0,1,1,1,1,0,0,1,1,0,0,0,0,1,1]
$\{I_{46}^{\text{PT2}},I_{47}^{\text{PT2}}\}$
Sector $\mathbf{F_{132}}$[1,1,1,0,0,0,0,1,1,0,0,1,1,0,1]
$\{I_{48}^{\text{PT3}},I_{49}^{\text{PT3}}\}$
Sector $\mathbf{F_{132}}$[0,1,0,0,0,0,1,1,1,0,0,1,1,1,1]
$\{I_{32}^{\text{PT3}},I_{33}^{\text{PT3}},I_{34}^{\text{PT3}},I_{35}^{\text{PT3}},I_{36}^{\text{PT3}}\}$
Sector $\mathbf{F_{123}}$[0,1,0,0,0,0,1,1,1,0,0,1,1,1,1]
$\{I_{37}^{\text{PT1}},I_{38}^{\text{PT1}},I_{39}^{\text{PT1}},I_{40}^{\text{PT1}},I_{41}^{\text{PT1}},I_{42}^{\text{PT1}},I_{43}^{\text{PT1}}\}$
Nine-Propagator Pure Candidates
Sector $\mathbf{F_{123}}$[0,1,1,1,1,0,0,1,1,1,0,0,0,1,1]
$\{I_{19}^{\text{PT2}},I_{20}^{\text{PT2}}\}$
Sector $\mathbf{F_{132}}$[1,1,1,0,0,0,1,1,1,0,0,1,1,0,1]
$\{I_{18}^{\text{PT3}},I_{19}^{\text{PT3}}\}$
Sector $\mathbf{F_{123}}$[1,1,1,0,0,0,1,1,1,0,0,1,0,1,1]
$\{I_{34}^{\text{PT1}},I_{35}^{\text{PT1}},I_{36}^{\text{PT1}}\}$
Sector $\mathbf{F_{123}}$[1,0,0,1,0,1,0,1,0,1,1,1,1,0,1]
$\{I_{7}^{\text{PL2}},I_{8}^{\text{PL2}},I_{9}^{\text{PL2}}\}$
Sector $\mathbf{F_{132}}$[0,1,1,0,1,0,1,0,1,0,1,1,1,0,1]
$\{I_{7}^{\text{PL3}},I_{8}^{\text{PL3}},I_{9}^{\text{PL3}}\}$
Sector $\mathbf{F_{123}}$[1,0,1,0,0,0,1,1,1,0,0,1,1,1,1]
$\{I_{27}^{\text{PT1}},I_{28}^{\text{PT1}},I_{29}^{\text{PT1}},I_{30}^{\text{PT1}}\}$
Sector $\mathbf{F_{132}}$[1,1,0,0,0,0,1,1,1,0,0,1,1,1,1]
$\{I_{14}^{\text{PT3}},I_{15}^{\text{PT3}},I_{16}^{\text{PT3}},I_{17}^{\text{PT3}}\}$
Sector $\mathbf{F_{132}}$[0,1,1,0,0,0,1,1,1,0,0,1,1,1,1]
$\begin{split}\{&I_{5}^{\text{PT3}},I_{6}^{\text{PT3}},I_{7}^{\text{PT3}},I_{8}^{\text{PT3}},I_{9}^{\text{PT3}},\\ &I_{10}^{\text{PT3}},I_{11}^{\text{PT3}},I_{12}^{\text{PT3}},I_{13}^{\text{PT3}}\}\end{split}$
Sector $\mathbf{F_{123}}$[1,1,0,0,0,0,1,1,1,0,0,1,1,1,1]
$\begin{split}\{&I_{17}^{\text{PT1}},I_{18}^{\text{PT1}},I_{19}^{\text{PT1}},I_{20}^{\text{PT1}},I_{21}^{\text{PT1}},\\ &I_{22}^{\text{PT1}},I_{23}^{\text{PT1}},I_{24}^{\text{PT1}},I_{25}^{\text{PT1}},I_{26}^{\text{PT1}}\}\end{split}$
Ten-Propagator Pure Candidates
Sector $\mathbf{F_{123}}$[1,1,0,1,0,1,0,1,0,1,1,1,1,0,1]
$\begin{split}\{&I_{1}^{\text{PL2}},I_{2}^{\text{PL2}},I_{3}^{\text{PL2}}\}\end{split}$
Sector $\mathbf{F_{132}}$[1,1,1,0,1,0,1,0,1,0,1,1,1,0,1]
$\begin{split}\{&I_{1}^{\text{PL3}},I_{2}^{\text{PL3}},I_{3}^{\text{PL3}}\}\end{split}$
Sector $\mathbf{F_{123}}$[1,1,1,0,0,0,1,1,1,0,0,1,1,1,1]
$\begin{split}\{&I_{1}^{\text{PT1}},I_{2}^{\text{PT1}},I_{3}^{\text{PT1}},I_{4}^{\text{PT1}},I_{5}^{\text{PT1}},I_{6}^{\text{PT1}}\}\end{split}$
Sector $\mathbf{F_{123}}$[0,1,1,1,1,0,0,1,1,1,0,0,1,1,1]
$\begin{split}\{&I_{1}^{\text{PT2}},I_{2}^{\text{PT2}},I_{3}^{\text{PT2}},I_{4}^{\text{PT2}}\}\end{split}$
Sector $\mathbf{F_{132}}$[01,1,1,0,0,0,1,1,1,0,0,1,1,1,1]
$\begin{split}\{&I_{1}^{\text{PT3}},I_{2}^{\text{PT3}},I_{3}^{\text{PT3}},I_{4}^{\text{PT3}}\}\end{split}$